OpenStudy (yanasidlinskiy):

. The volume in cubic feet of a box can be expressed as (x)=x^3-〖6x〗^2+8x , or as the product of three linear factors with integer coefficients. The width of the box is 2 – x. Factor the polynomial to find linear expressions for the height and the length

jimthompson5910 (jim_thompson5910):

what is the GCF of: x^3, -6x^2, 8x

jimthompson5910 (jim_thompson5910):

wait, is it (6x)^2 or 6x^2 ?

OpenStudy (yanasidlinskiy):

Let me check.

OpenStudy (yanasidlinskiy):

Sorry..\[-6x^2\]

jimthompson5910 (jim_thompson5910):

that's fine

jimthompson5910 (jim_thompson5910):

what is the GCF of: x^3, -6x^2, 8x ?

OpenStudy (yanasidlinskiy):

Anyways...Would it be 2?

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (yanasidlinskiy):

I don't know then..I was thinking 2. The entire time.

OpenStudy (yanasidlinskiy):

How come it's not 2?

jimthompson5910 (jim_thompson5910):

because 2 isn't found in the coefficient of x^3

jimthompson5910 (jim_thompson5910):

what factor is in all three terms: x^3, -6x^2, 8x ?

OpenStudy (yanasidlinskiy):

Oh..Ok. x? I really don't knowxD

jimthompson5910 (jim_thompson5910):

yep x

OpenStudy (yanasidlinskiy):

Hahha:D Wow.

jimthompson5910 (jim_thompson5910):

that means x^3-6x^2+8x = x(x^2 - 6x + 8)

jimthompson5910 (jim_thompson5910):

now factor x^2 - 6x + 8 to get ???

OpenStudy (yanasidlinskiy):

Umm. x^2 - 6x + 8 = x(x^2 - 6x + 8). Right?

jimthompson5910 (jim_thompson5910):

To factor x^2 - 6x + 8, you need to find two numbers that multiply to 8 (third term) AND add to -6 (second coefficient)

OpenStudy (yanasidlinskiy):

Ok. Can you show me step by step how? I just got into this class and I don't have a clue.

OpenStudy (yanasidlinskiy):

Oh, wait. Would it be 2?

jimthompson5910 (jim_thompson5910):

what are some ways to factor 8?

OpenStudy (yanasidlinskiy):

Well 2*4 = 8.

OpenStudy (yanasidlinskiy):

(2*-4)(2*4) <---Is that right?

jimthompson5910 (jim_thompson5910):

notice how 2+4 = 6

jimthompson5910 (jim_thompson5910):

we want the two factors to add to -6, not +6

jimthompson5910 (jim_thompson5910):

we can easily fix this by saying (-2)*(-4) = 8 -2 + (-4) = -6

jimthompson5910 (jim_thompson5910):

therefore, x^2 - 6x + 8 = (x-2)(x-4)

OpenStudy (yanasidlinskiy):

And that would be my final answer?

jimthompson5910 (jim_thompson5910):

so we have this full factorization x^3-6x^2+8x = x(x-2)(x-4)

jimthompson5910 (jim_thompson5910):

it says "The width of the box is 2 – x. " notice how we have x-2 and not 2-x as a factor

jimthompson5910 (jim_thompson5910):

we can easily fix this by pulling out a negative from each x-2 and x-4 x-2 = -1(-x+2) = -1(2-x) x-4 = -1(-x+4) = -1(4-x)

jimthompson5910 (jim_thompson5910):

so, x^3-6x^2+8x = x(x-2)(x-4) x^3-6x^2+8x = x( -1(2-x) )( -1(4-x) ) x^3-6x^2+8x = x*( -1)*(-1)(2-x)(4-x) x^3-6x^2+8x = x*(1)(2-x)(4-x) x^3-6x^2+8x = x(2-x)(4-x)

OpenStudy (yanasidlinskiy):

Wow! That's a lot. And..That makes more sense. Would that be all?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

so that means the product of the length and height is x(4-x)

jimthompson5910 (jim_thompson5910):

the length could be x, or it could be 4-x the height could be x, or it could be 4-x

jimthompson5910 (jim_thompson5910):

it really doesn't matter since you can multiply in any order

OpenStudy (yanasidlinskiy):

Oh, wow. My teacher like "preached" to me that It does matter. She made a lot of mistakes anyways. So I'll rather trust you more than her:D

jimthompson5910 (jim_thompson5910):

here is an example 2*4 = 8 4*2 = 8

jimthompson5910 (jim_thompson5910):

multiplying can be done in any order

OpenStudy (yanasidlinskiy):

Exactly!!:D I tried to tell her and she's like..No, no no. You're wrong. And I'm like...Whatever.

jimthompson5910 (jim_thompson5910):

strange, that's basic math that any teacher would know

OpenStudy (yanasidlinskiy):

Hahha:D Yep! I don't even know how she got hired. Lol.

jimthompson5910 (jim_thompson5910):

A strange reality in education, yeah I know